1. Assume that Roberts’s utility from consuming good X and good Y is given by the following function:


U = X0.3Y0.7


Where X is the quantity of good X while Y is the quantity of good Y.


Assume the price of X (PX) is £25, the price of Y (PY) is £35 and he has a budget of £1000 to spend on the two goods. 



c.     Using the demand functions, calculate the quantities of X and Y Robert should purchase to maximise his utility. Calculate the utility this optimal consumption bundle provides.






1
Expert's answer
2020-11-16T12:00:35-0500

Objective: Maximise "X^{0.3}Y^{0.7}"

Subject to:

25X + 35Y = 1000

x≥ 0, y≥0

Px = £25

Py = £35

TC = Px *X + Py*Y

"TC = P_{x} \times X + P_{y} \times Y"

TC = 25X + 35Y = 1000

f(x,y) = "X^{0.3}Y^{0.7}"

g(x,y) = 25X + 35Y = 1000

L(λ, x, y) = X0.3Y0.7 − λ(25X + 35Y - 1000)

0.3X-0.7Y0.7 = 25λ

0.7X0.3Y-0.3 = 35λ

25X + 35Y = 1000

"\frac{0.3}{0.7}X^{-0.7-0.3}Y^{0.7 – - 0.3}" = "\frac{25λ}{35λ}"

"0.428571X^{-1}Y^{1}" = 0.714286

Y = 1.666669X

25X + 35Y = 1000

25X + 35(1.666669X) = 1000

25X + 58.33342X = 1000

83.33342X = 1000

X = "\frac{1000}{83.33342}"

X = 12

Y = 20

U = "X^{0.3}Y^{0.7}"

U = "12^{0.3}20^{0.7}"

U = "2.107436 \times 8.141811"

U = 17.15834